# Stochastic analysis of finite-temperature effects on cosmological parameters by artificial neural networks

### v1

@@ Abstract
We explore the impact of finite-temperature quantum gravity effects on cosmological parameters, particularly the cosmological constant $\Lambda$, by incorporating temperature-dependent quantum corrections into the Hubble parameter. For that purpose, we modify the Cosmic Linear Anisotropy Solving System. We introduce new density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, arising from finite-temperature quantum gravity contributions, and analyze their influence on the cosmic microwave background power spectrum using advanced machine learning techniques, including artificial neural networks and stochastic optimization. Our results reveal that $\Omega_{\Lambda_{3}}$ assumes a negative value, consistent with dimensional regularization in renormalization and that the presence of $\Omega_{\Lambda_{2}}$ as well as $\Omega_{\Lambda_{3}}$ significantly enhances model accuracy. Numerical analyses demonstrate that the inclusion of these parameters improves the fit to 2018 Planck data, suggesting that finite-temperature quantum gravity effects play a non-negligible role in cosmological evolution. Although the Hubble tension persists, our findings highlight the potential of quantum gravitational corrections in refining cosmological models and motivate further investigation into higher-order thermal effects and polarization data constraints.
@@ 1 Introduction
The cosmological constant problem remains one of the most profound challenges in theoretical cosmology (see [1, 2] for reviews), with the observed value of vacuum energy being many orders of magnitude smaller than quantum field theory predictions. Recent works [3, 4, 5] employing finite-temperature quantum field theory (finite-T QFT) (see [6, 7, 8] for reviews) suggest that temperature-dependent quantum corrections may provide new insights into this longstanding puzzle. These effects, particularly in the context of quantum gravity (QG), introduce perturbative corrections to the cosmological constant that could significantly impact our understanding of early Universe physics.
@@
In conventional cosmology [9, 10], temperature effects are typically treated through classical thermodynamics, while quantum contributions are disregarded. However, as demonstrated in [3, 4], where it was proposed that the cosmological constant (CC) problem may be a consequence of how perturbation theory and finite renormalization are handled (see [11] and [12] for related discussions), finite-T QFT effects naturally lead to a time-dependent cosmological constant in FLRW cosmology through loop-generated corrections. This arises from the temperature dependence of the one-particle irreducible (1PI) effective action, where the quantum-corrected cosmological constant takes the form $\Lambda_{tot}=\Lambda_{1}+\Lambda_{2}a^{-4}+\Lambda_{3}a^{-2}+\cdots$ with $a$ the scale factor, introducing new density parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ beyond the standard $\Omega_{\Lambda_{1}}$.
@@
This work presents a comprehensive investigation of these finite-temperature QG effects by conducting the following tasks: firstly, we modify the Cosmic Linear Anisotropy Solving System (CLASS) Boltzmann code [13, 14] to incorporate temperature-dependent corrections to the Hubble parameter, including the new QG-induced parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$. Secondly, we carry out an extensive numerical analysis by using both brute-force parameter scans and advanced machine learning techniques to constrain these parameters against the 2018 Planck data [15]. Thirdly, we develop artificial neural network (ANN) methods to efficiently explore the high-dimensional parameter space and identify optimal cosmological configurations
@@
The inclusion of finite-T QG effects addresses several key theoretical questions. First, the negative sign that emerges for $\Omega_{\Lambda_{2}}$ in our analysis finds natural explanation through dimensional regularization in the renormalization procedure. Second, while these effects are small in the present era, they become significant during the radiation-dominated epoch and near recombination, potentially influencing the interpretation of spatial curvature and early dark energy [16]. Third, our phenomenological approach provides a framework to test whether quantum gravitational corrections could alleviate persistent tensions like the Hubble tension [17].
@@
Our methodology combines first-principles theoretical considerations with state-of-the-art computational techniques. The modified CLASS implementation [18] allows us to systematically study how finite-T corrections affect the CMB power spectrum, while our machine learning (ML) approach enables efficient exploration of the eight-dimensional parameter space ($\Omega_{\Lambda_{2}}$, $\Omega_{\Lambda_{3}}$, $h$, $\omega_{b}$, $\omega_{cdm}$, $A_{s}$, $n_{s}$, $\tau_{reio}$). Feature importance analysis reveals that $\Omega_{\Lambda_{2}}$ plays a more significant role than several established parameters in predicting cosmological observables.
The paper is organized as follows: section 2 details our modifications to CLASS and provides a review of the statistical and ML techniques employed in the main analysis. Section 3 presents our numerical methods and results, including ANN-based parameter estimation. We conclude with discussion of implications for cosmological models and future directions.
2 Computational frameworks
@@
In this section, we outline the two main computational tools developed and employed in this study to incorporate finite-T effects and analyze their cosmological implications. First, we describe the modifications made to the CLASS Boltzmann code to account for finite-temperature quantum gravitational corrections, which manifest as shifts in cosmological parameters, most notably the cosmological constant. These modifications enable the computation of power spectra under the influence of temperature-dependent vacuum energy contributions. Second, we present a statistical framework based on ANNs. This ANN-based model is further used for sensitivity analysis through feature ablation, as well as for probabilistic optimization of cosmological parameters.
{fmffile}
[…]
@@
Before delving into the specific computational tools used in this work, it is useful to outline a more conceptual framework that motivates our overall approach. At a fundamental level, the starting point would be the action describing the standard model (SM) fields coupled to gravity. To this, one may append the action of an additional macroscopic system, such as a hydrodynamic fluid, which serves to model the large-scale structure and thermodynamic evolution of the Universe. Given that quantum effects are expected to play a significant role - particularly in the early Universe — it is natural to consider the 1PI effective action for the entire coupled system. More specifically, one can take the classical action of the form
@@
where $S_{EH}$ denotes an Einstein-Hilbert action with a cosmological constant, $S_{SM}$ the SM action (in a curved background), and $S_{PF}$ a hydrodynamic matter system, say, a perfect fluid, to be specific:
@@
where the contractions of the spacetime indices are done with the curved metric; the covariant derivatives contain both the Christoffel and gauge connections. The detailed forms and field contents of each line above as well as the overall conventions can be found in [5]. The explicit form of the action $S_{PF}$ can be found, e.g., in[19]. One can quantize the system [5], and obtain the 1PI action. For cosmological applications, the important terms will be the classical part and leading terms in the derivative expansion, as higher-order contributions are typically suppressed. More precisely, one can focus on the renormalized action with the coupling constants shifted by quantum effects.
@@
Especially, non-derivative quantum corrections, such as those contributing to the cosmological constant, can have a non-negligible impact despite their small magnitude. This is so because the observed value of the cosmological constant is itself extremely small, making even suppressed quantum effects potentially significant. Similar roles for non-derivative quantum corrections have been explored in other contexts, such as black hole accretion physics [20], underscoring their physical relevance. Shifts in the cosmological constant arise from the vacuum-to-vacuum diagrams. Fig. 1 shows the relevant diagrams in the pure gravity sector (analogous diagrams exist in the matter-coupled sector). For example, the graviton vacuum-to-vacuum amplitude is given by:
@@
$\int\prod_{x}dh_{\kappa_{1}\kappa_{2}};e^{\frac{i}{\kappa^{\prime 2}}\int\sqrt{-\tilde{g}},\left(-\frac{1}{2}\tilde{\nabla}\gamma h^{\alpha\beta}\tilde{\nabla}^{\gamma}h{\alpha\beta}\right)}.$ (3)
@@
Evaluation of such diagrams in a finite temperature generically leads to the form of the CC given in the introduction.

### v2

@@ Abstract
We explore the impact of finite-temperature quantum gravity effects on cosmological parameters, particularly the cosmological constant $\Lambda$, by incorporating temperature-dependent quantum corrections into the Hubble parameter. For that purpose, we modify the Cosmic Linear Anisotropy Solving System. We introduce new density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, arising from finite-temperature quantum gravity contributions, and analyze their influence on the cosmic microwave background power spectrum using advanced machine learning techniques, including artificial neural networks and stochastic optimization. Our results reveal that $\Omega_{\Lambda_{2}}$ assumes a negative value, consistent with dimensional regularization in renormalization and that the presence of $\Omega_{\Lambda_{2}}$ as well as $\Omega_{\Lambda_{3}}$ enhances model accuracy. Numerical analyses demonstrate that the inclusion of these parameters improves the fit to 2018 Planck data. Although further work is required, our results suggest that finite-temperature quantum gravity effects may play a non-negligible role in cosmological evolution. Although the Hubble tension persists, our findings highlight the potential of quantum gravitational corrections in refining cosmological models and motivate further investigation into higher-order thermal effects and polarization data constraints.
@@ 1 Introduction
The cosmological constant problem remains one of the most profound challenges in theoretical cosmology (see [1, 2, 3] for reviews and recent developments), with the observed value of vacuum energy being many orders of magnitude smaller than quantum field theory predictions. Recent works [4, 5, 6] employing finite-temperature quantum field theory (finite-T QFT) (see [7, 8, 9] for reviews) suggest that temperature-dependent quantum corrections may provide new insights into this longstanding puzzle. [^4] The central idea in [4, 5] is the introduction of a novel renormalization scheme in which the renormalized masses are chosen to be of the order of the temperature. Similar approaches were adopted in earlier works, such as [10], where problematic $m^{4}$ terms were eliminated through an alternative renormalization prescription. These effects, particularly in the context of quantum gravity (QG), introduce perturbative corrections to the cosmological constant that could significantly impact our understanding of early Universe physics.
@@
In conventional cosmology (see [11, 12] for reviews), temperature effects are typically treated through classical thermodynamics, while quantum contributions are disregarded. However, as demonstrated in [4, 5], where it was proposed that the cosmological constant (CC) problem may be a consequence of how perturbation theory and finite renormalization are handled (see [1, 2], [13], and [14] for related discussions), finite-T QFT effects naturally lead to a time-dependent cosmological constant in FLRW cosmology through loop-generated corrections. This arises from the temperature dependence of the one-particle irreducible (1PI) effective action, where the quantum-corrected cosmological constant takes the form $\Lambda_{tot}=\Lambda_{1}+\Lambda_{2}a^{-4}+\Lambda_{3}a^{-2}+\cdots$ with $a$ the scale factor, introducing new density parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ beyond the standard $\Omega_{\Lambda_{1}}$.
@@
Continuing the effort initiated in [15], this work presents an intensive investigation of these finite-temperature QG effects on the cosmological parameters by conducting the following tasks: firstly, we modify the Cosmic Linear Anisotropy Solving System (CLASS) Boltzmann code [16, 17] to incorporate temperature-dependent corrections to the Hubble parameter, including the new QG-induced parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$. Secondly, we carry out an extensive numerical analysis by using both brute-force parameter scans and advanced machine learning techniques to constrain these parameters against the 2018 Planck data [18]. Thirdly, we develop artificial neural network (ANN) methods to efficiently explore the high-dimensional parameter space and identify optimal cosmological configurations. Studies employing ANNs to investigate cosmological parameters include [21, 19, 22, 20, 23].
@@
The inclusion of finite-T QG effects addresses several key theoretical questions. First, the negative sign that emerges for $\Omega_{\Lambda_{2}}$ in our analysis finds natural explanation through dimensional regularization in the renormalization procedure. Second, while these effects are small in the present era, they become significant during the radiation-dominated epoch and near recombination, potentially influencing the interpretation of early dark energy [24]. Third, our phenomenological approach provides a framework to test whether quantum gravitational corrections could alleviate persistent tensions like the Hubble tension [25].
@@
Our methodology combines first-principle theoretical considerations with state-of-the-art computational techniques. The modified CLASS implementation [26] allows us to systematically study how finite-T corrections affect the CMB power spectrum, while our machine learning (ML) approach enables efficient exploration of the eight-dimensional parameter space ($\Omega_{\Lambda_{2}}$, $\Omega_{\Lambda_{3}}$, $h$, $\omega_{b}$, $\omega_{cdm}$, $A_{s}$, $n_{s}$, $\tau_{reio}$).
The paper is organized as follows. Section 2 describes our modifications to CLASS. A natural question arises: given their similar scaling behaviors, could $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ be absorbed into existing parameters such as the radiation density $\Omega_{r}$ and the curvature density $\Omega_{K}$, respectively? We address this question. Section 3 outlines our numerical methods and results, including parameter estimation using artificial neural networks (ANNs). Our approach, which is a model-building exercise, differs from the standard Bayesian parameter inference, adopting a frequentist framework augmented by machine learning techniques; the rationale behind this choice is also discussed. It is noteworthy that both the results obtained in Table 3 and Fig. 7 indicate that $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ play a crucial role in determining the fine structure of the curve. In section 4, we conclude with a discussion of the implications for cosmological models and potential future directions. In Appendix A, the non-degeneracy of $\Omega_{\Lambda_{3}}$ and $\Omega_{K}$ is explicitly demonstrated. Appendix B presents the ANN-based statistical method employed in the main analysis of section 3.
2 Computational framework
@@
In this section, we outline the main computational tool developed and employed in this study to incorporate finite-T effects and analyze their cosmological implications, a modified version of CLASS Boltzmann code. We describe the modifications made to the CLASS to account for finite-temperature quantum gravitational corrections, which manifest as shifts in cosmological parameters, most notably the cosmological constant. These modifications enable the computation of power spectra under the influence of temperature-dependent vacuum energy contributions.
2.1 Conceptual framework
@@
Before delving into the specific computational tools used in this work, it is useful to outline a more conceptual framework that motivates our overall approach. At a fundamental level, the starting point would be the action describing the Standard Model (SM) fields coupled to gravity. To this, one may append the action of an additional macroscopic system, such as a hydrodynamic fluid, which serves to model the large-scale structure and thermodynamic evolution of the Universe. Given that quantum effects are expected to play a significant role - particularly in the early Universe - it is natural to consider the 1PI effective action for the entire coupled system. More specifically, one can take the classical action of the form
@@
where $S_{EH}$ denotes an Einstein-Hilbert action with a cosmological constant, $S_{SM}$ the SM action (in a curved background), and $S_{PF}$ a hydrodynamic or kinetic theory matter system, say, a perfect fluid, to be specific:
@@
$-Y_{mn}^{d}\,\bar{q}_{L}^{m}\Phi\,{d_{R}}^{n}-Y_{mn}^{u}\,\bar{q}_{L}^{m}\tilde{{\cal F}}\,u_{R}^{n}+h.c.+\mbox{(leptonic sector Yukawa terms)}$
where the contractions of the spacetime indices are done with the curved metric; the covariant derivatives contain both the Christoffel and gauge connections. The detailed forms and field contents of each line above as well as the overall conventions can be found in [6]. The explicit form of the action $S_{PF}$ in the hydrodynamic matter case can be found, e.g., in[27]. One can quantize the system [6], and obtain the 1PI action. For cosmological applications, the important terms will be the classical part (or renormalized part, more precisely) and leading terms in the derivative expansion, as higher-order contributions are typically suppressed. More precisely, one can focus on the renormalized action with the coupling constants shifted by quantum effects. [^5] As a slightly different and conceptually simpler approach - yet one that ultimately leads to the same starting point at the level of approximation currently adopted for cosmological studies as just described - one may employ the following setup. In this approach, one first constructs an effective field theory action without incorporating the kinetic theory component. The term ”effective” can carry different meanings in the literature; here, it refers specifically to an effective field theory obtained by retaining only the leading terms in the derivative expansion of the 1PI action. Notably, quantum corrections to the cosmological constant are included. The kinetic theory system can then be coupled to this resulting field-theoretic system.
{fmffile}
@@
(a)            (b)
@@
Especially, non-derivative quantum corrections, such as those contributing to the cosmological constant, can have a non-negligible impact despite their small magnitude. This is so because the observed value of the cosmological constant is itself extremely small, making even suppressed quantum effects potentially significant. Similar roles for non-derivative quantum corrections have been explored in other contexts, such as black hole accretion physics [28], underscoring their physical relevance in some situations. Shifts in the cosmological constant arise from the vacuum-to-vacuum diagrams. Fig. 1 shows the relevant diagrams in the pure gravity sector (analogous diagrams exist in the matter-coupled sector). For example, the graviton vacuum-to-vacuum amplitude is given by:
@@
$\int\prod_{x}dh_{\kappa_{1}\kappa_{2}}\;e^{\frac{i}{\kappa^{\prime 2}}\int\sqrt{-\tilde{g}}\,\left(-\frac{1}{2}\tilde{\nabla}\gamma h^{\alpha\beta}\tilde{\nabla}^{\gamma}h_{\alpha\beta}\right)}.$ (3)
@@
Evaluation of such diagrams including the ones coming from the SM sector in a finite temperature, an example of which is given below, generically leads to the form of the CC given in the introduction [4, 5, 6]. For relavance of dynamic vacuum in the context of the CC problem, see [29, 30]
To avoid unnecessary complications associated with gravity quantization, for the moment let us consider s scalar system to illustrate one-loop generation of the CC. In a flat background:
$S_{ren}$ $=$ $-\int d^{4}x\;\left[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4!}{\lambda}\phi^{4}\right].$ (4)
where ’ren’ denotes ’renormalized’. Define
$V_{ren}\equiv\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4!}{\lambda}\phi^{4}\qquad,\qquad m_{\phi}\equiv m^{2}+\frac{1}{2}\phi^{2}.$ (5)
The one-loop effective potential at finite temperature is given by (see e.g., [31], the work well-known in the finite-T literature, as well as [9])
$V_{one-loop}=V_{ren}+J(m_{\phi},T)$ (6)
$J(m_{\phi},T)$ $\equiv$ $\;\;\mathclap{\displaystyle\int}\mathclap{\textstyle\sum}\;\;\;\frac{1}{2}\ln(K^{2}+m_{\phi}^{2})=-\frac{\pi^{2}}{90}\,T^{4}+\frac{m_{\phi}^{2}}{24}\,T^{2}-\frac{1}{12\pi}m_{\phi}^{3}\,T$ (7)
$\hskip-50.58878pt-\frac{\mu^{-2\epsilon}}{2(4\pi)^{2}}\;m_{\phi}^{4}\Big[\ln\Big(\frac{\bar{\mu}e^{\gamma}}{4\pi T}\Big.)+\frac{1}{2\epsilon}\Big]+\frac{\zeta(3)}{3(4\pi)^{4}}\frac{m_{\phi}^{6}}{T^{2}}+{\cal O}\Big(\frac{m_{\phi}^{8}}{T^{4}}\Big)+{\cal O}(\epsilon)$
where $\ln\bar{\mu}^{2}\equiv\ln(4\pi\mu^{2})-\gamma$ with $\mu$ being the scale parameter of dimensional regularization; $\gamma$ is an Euler’s constant; $\epsilon$ is a regularization parameter of dimensional regularization; $\zeta(3)$ is zeta function evaluated at 3. For the contributions to the CC, one can note the $\phi$-independent terms: the first two leading order terms scale $\sim T^{4}$ and $\sim T^{2}$. To extend the discussion to an FLRW background, let us extend the system to a curved background.
$S_{ren}=\frac{1}{\kappa^{2}}\int d^{4}x\sqrt{\gamma}\;(R-2\Lambda_{1})-\int d^{4}x\sqrt{-g}\;\Big(\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+V\Big)$ (8)
where $\kappa^{2}=16\pi G$ with $G$ Newton’s constant; $\kappa^{2}$ will be suppressed in some places. As well known, an FLRW background has a time-dependent temperature. Due to this, the standard finite-temperature cannot in general be directly applied. This issue has been recently addressed in [6]. For an FLRW background, the standard finite-T techniques can be applied with minor modifications. This is due to the special feature of an FLRW background (the flat case in which we are interested) that the temperature scales as $\sim\frac{1}{a}$ and the scale factor happens to be a metric component:
$ds^{2}=-dt^{2}+a(dx^{2}+dy^{2}+dz^{2})$ (9)
Therefore the temperature is annihilated by the covariant derivative and the upshot is that the temperature $T$ appearing in the results such as (17) can simply be replaced by the time-dependent one $T(t)$. The result (17) thus implies
$\Lambda\equiv\Lambda_{1}+\frac{\Lambda_{2}}{a^{4}}+\frac{\Lambda_{3}}{a^{2}}+\cdots$ (10)
where $\Lambda_{2}$ and $\Lambda_{3}$ are numerical constant introduced as part of the definition of $\Lambda.$ $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ are then introduced as the corresponding density parameters, which implies the modified form of the Hubble constant given in (21).
[…]
${\cal L}=\frac{1}{\kappa}(-2\Lambda_{1}+R)$ (11)
We expand the metric $g_{\mu\nu}$ around a background solution $g_{0\mu\nu}(K)$ such that $g_{\mu\nu}\equiv h_{\mu\nu}+g_{0\mu\nu}(K)$, where the background is given by:
$ds^{2}=a^{2}\left[d{\bf x}^{2}+K\frac{({\bf x}\cdot d{\bf x})^{2}}{1-K{\bf x}^{2}}\right]$ (12)
By substituting this expansion into the action and taking the $h_{\mu\nu}$-variation, one obtains the field equations. In this context, the background curvature parameter $K$ appears in multiple locations within the resulting equations. This explains why the density parameter $\Omega_{K}$, defined as
$\Omega_{K}\equiv-\frac{K}{a_{0}^{2}H_{0}^{2}}$ (13)
enters the standard cosmological equations through several distinct channels. In contrast, the parameter $\Omega_{\Lambda_{3}}$ has an entirely different physical origin. When finite-temperature one-loop corrections are incorporated, the cosmological constant $\Lambda_{1}$ is shifted schematically as:
$\Lambda_{1}+\Lambda_{2}\frac{1}{a^{4}}+\Lambda_{3}\frac{1}{a^{2}}+\cdots$ (14)
Unlike $\Omega_{K}$, the parameter $\Omega_{\Lambda_{3}}$ originates exclusively from the cosmological-constant term of the off-shell one-loop effective action. Following a standard linearized analysis of gravity coupled with hydrodynamic matter, it enters the Hubble parameter HH. [^6] The fluctuations $\delta\rho_{\Lambda_{2}}$ and $\delta\rho_{\Lambda_{3}}$ are neglected due to their suppressed nature - being both quantum in origin and small variations. Thus, their influence is minimal, and $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ effectively appear only within the modified Hubble parameter $H$. (This statement holds at the level of approximation adopted in the present work. At higher orders, additional terms such as $\Lambda R_{\mu\nu}R^{\mu\nu}$ are expected to appear.) In contrast, $\Omega_{K}$ and $\Omega_{r}$ directly contribute to the fluctuation equations and plays a significant role in the evolution of perturbations. In this work, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ are treated phenomenologically, and their inclusion is justified by consistency with numerical results. (Theoretically, these parameters can be computed from first principles via the renormalization procedure. However, finite renormalization freedom necessitates experimental inputs like the present one to fix their values.) While $\Omega_{K}$ (once kept) and $\Omega_{\Lambda_{3}}$ appear only in the combination $\Omega_{\Lambda_{3}}+\Omega_{K}$ within the expression for $H$, these parameters are fundamentally non-redundant
  • Origins: One is geometric (curvature), while the other stems from quantum corrections (loop effects).
  • Appearances: They enter the action and resulting field equations through distinct mathematical terms.
The non-degeneracy of these parameters has been explicitly verified in Appendix A through numerical analysis using a modified version of the CLASS code.
As for $\Omega_{\Lambda_{2}}$, one might also think that finite-temperature quantum loop-induced contributions to the CC - especially those involving photons and neutrinos - are redundant, since these are already accounted for as radiation. However, this conflation is misleading. Radiation arises from physical, on-shell particles (e.g., photons, neutrinos) through, say, kinetic theory physics, whereas the contributions under consideration are due to virtual particles in loop corrections. The finite-temperature effects we consider are analogous to vacuum energy contributions induced by loop effects in QED photon and electron fields. Radiation density, on the other hand, pertains to the energy density of physical photons and neutrinos. Therefore, the finite-temperature effects we consider are not included in the conventional radiation density and must be accounted for separately.
To be precise, let us examine a system governed by Boltzmann kinetic theory - such as a photon gas - interacting with a scalar field (e.g., the Higgs field) in a curved spacetime background. The total action of the system is given by
$S=S_{EH}+S_{sf}+S_{kin},$ (15)
where $S_{EH}$ represents the Einstein-Hilbert action with a cosmological constant, $S_{sf}$ is the scalar field action, and $S_{kin}$ accounts for the kinetic sector. While the explicit form of $S_{kin}$ is known in certain models (see, e.g., [32, 33]) - such as that of an ideal gas - it will not be needed for our current purposes. The associated metric field equation contains the stress-energy tensor:
$T^{\mu\nu}=\frac{1}{\sqrt{-g}}\sum_{n}\delta^{3}(x_{n}(t)-x)\,p_{n}^{\mu}(t)\,p_{n}^{\nu}(t)\,E_{n}(t).$ (16)
To avoid unnecessary complications, we assume a fixed, non-dynamical background metric. Our primary interest lies in the $T^{4}$-scaling (where T denotes the temperature) behavior relevant to $\Omega_{\Lambda_{2}}$. The conventional analysis focuses exclusively on the classical thermodynamics of the kinetic system, which leads to the well-known exact $T^{4}$-dependence of the energy density - see, for example, footnote 1 in section 3.1 of [11]. In contrast, the focus of this work is on quantum field-theoretic effects originating from the scalar field sector $S_{sf}$ (as well as from other Standard Model particles and the graviton in the full description, which do not qualitatively alter the conclusions) at finite temperature. Importantly, such scalar fields cannot be treated as radiation due to their large masses; instead, their contributions to the cosmological constant arise through virtual processes, specifically vacuum bubble diagrams. To be concrete, consider the scalar sector. The one-loop correction to the classical (more precisely, renormalized) action takes the form
$\mathclap{\displaystyle\int}\mathclap{\textstyle\sum}\;\;\;\frac{1}{2}\ln(K^{2}+m^{2})=-\frac{\pi^{2}}{90}\,T^{4}+\frac{m^{2}}{24}\,T^{2}-\frac{1}{12\pi}m^{3}\,T$
$\hskip-36.135pt-\frac{\mu^{-2\epsilon}}{2(4\pi)^{2}}\;m^{4}\Big[\ln\Big(\frac{\bar{\mu}e^{\gamma}}{4\pi T}\Big.)+\frac{1}{2\epsilon}\Big]+\frac{\zeta(3)}{3(4\pi)^{4}}\frac{m^{6}}{T^{2}}+{\cal O}\Big(\frac{m^{8}}{T^{4}}\Big)+{\cal O}(\epsilon)$ (17)
where $\ln\bar{\mu}^{2}\equiv\ln(4\pi\mu^{2})-\gamma$; $\gamma$ is an Euler’s constant; $\epsilon\equiv\frac{4-D}{2}$. This expression represents an infinite series expansion in powers of temperature, with the leading contribution scaling as $T^{4}$. Importantly, this term explicitly shifts the classical cosmological constant (CC) and is distinct from the $T^{4}$ contribution arising from the kinetic sector discussed earlier.